A company recently paid a dividend of $4.00 (Do). The dividend has been growing at a rate of 5% and this growth rate will continue until the end of year 5 and then starting with year 6, the growth rate of dividends will fall to 3% indefinitely. What is the Present Value of the dividends' cash stream to the investor starting with the end of year 1 based on theses projections if the required rate of return is 8%?
Answer #1

Step 1: Dividend to be payable in future Do Given $4.00 D1 = $4 + 5% $4.20 D2 =$4.20+5% $4.41 D3 =$4.41+5% $4.63 D4 =$4.63+5%Step 2: Since investor starting with the end of year 1, hence do not consider the dividend payable in year 1 i.e D1. Present